# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v2

@@ Appendix D Designed Prompts for Experiments
Appendix E Reflection Reports for Text Refinement
@@
Figure 20: Reflection summary for text refinement of Weather dataset.
Figure 21: Reflection summary for text refinement of Finance dataset.
Figure 22: Reflection summary for text refinement of Healthcare (Test-Positive) dataset.
Figure 23: Reflection summary for text refinement of Healthcare (Mortality) dataset.
Appendix F TimeXL for Regression-based Prediction: A Finance Demonstration
@@
We conduct the experiments on the same finance dataset, where the target value is the raw material stock price. The performance of all baselines and TimeXL is presented in Table 6. Consistent with the classification setting, TimeXL achieves the best results, and the multi-modal variants of state-of-the-art baselines (MM-iTransformer and MM-PatchTST) benefit from incorporating additional text data. This observation is further supported by the ablation study in Table 7. We also note that the text modality offers complementary information for LLMs, enhancing numerical price predictions. Furthermore, the identified prototypes provide contextual guidance, leading to additional performance gains. Overall, TimeXL outperforms all variants, underscoring the effectiveness of mutually augmented prediction. We also provide an iteration analysis to show the effectiveness of reflective and refinement process, as shown in Table 8. The prediction performance quickly improves and stabilizes over iterations, which underscores the alternation steps between predictions and reflective refinements. Additionally, we provide a visualization in Figure 26, which compares the predicted trend and stock price values, with ground truths for reference.
@@
Figure 26: Visualization of TimeXL’s numerical and trend predictions on Finance dataset.
Finally, we provide a case study on the testing set of finance data that demonstrate the effectiveness of TimeXL’s reasoning process for regression tasks. We present the top three most similar prototypes allocated for both time series and text modalities, where we have two key observations. 1) Even if the trend predictions align correctly in both original and refined texts, the prototype allocation varies regarding market sentiments. In the original texts, prototypes from the neutral class are also allocated, highlighting a mixed sentiment and potential fluctuations, consequently deteriorating numerical prediction accuracy from both the encoder and the LLM. In the refined texts, prototypes instead capture the underlying bearish market sentiment for iron ore, emphasizing decreased demand and weakened market conditions, effectively aligning bearish market sentiment with numerical predictions. 2) The allocated top similar time series prototypes also exhibit highly similar trends compared to the original input series, offering robust reference points for numerical prediction. To quantitatively assess it, we aggregate the retrieved prototypes using their similarity scores and compute the correlation between the input time series and the aggregated prototype series. The observed strong correlations underscores TimeXL’s efficacy in capturing intrinsic temporal patterns and providing predictive insights into future numerical stock values based on similar historical patterns.
Figure 27: Multi-modal case-based reasoning example on Finance dataset for regression tasks.
Appendix G More Detailed Discussions of Related Work
@@
G.1 Multi-modal Time Series Analysis.
@@
In recent years, multi-modal time series analysis has gained significant traction in diverse domains such as finance, healthcare, environmental sciences, and industry [11, 38, 12, 39, 16, 14]. Multiple approaches have been proposed to model interactions across different modalities, such as multi-modal fusion and cross-modal alignment and so on. Multi-modal fusion captures complementary information to enhance time series modeling, which are commonly implemented via addition [47] and concatenation [40, 70, 47, 41] at the multiple levels. For example, Liu et al. [47] fuse predictions from both time series and text modalities via learnable linear weighted addition, further enhancing the predictive performance of state-of-the-art forecasters. In parallel, cross-modal alignment aims to capture the association between time series and other modalities for downstream tasks, where common techniques include attention mechanism [13, 42, 104, 43], gating functions [44], and contrastive learning [45, 46]. For instance, Bamford et al. [17] align time series and text in a shared latent space using deep encoders, enabling retrieval of specific time series sequences based on textual queries. In addition, Zheng et al. [45] perform causal structure learning in multi-modal time series by separating modality-invariant and modality-specific components through contrastive learning, enabling root cause analysis. Besides method developments, recent research efforts have also advanced the field through comprehensive benchmarks and studies [47, 48, 49, 50, 51], highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore representing time series as other modalities (e.g., image, text, table, graph) [52, 53, 54, 55, 105] or generating synthetic modalities [44, 56] to facilitate downstream tasks. Nevertheless, these studies tend to focus primarily on improving numerical accuracy. The deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
G.2 Time Series Explanation
@@
Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based explanations leverage saliency methods to highlight important features at different time steps [57, 58], where some methods explicitly incorporate temporal structures into models and objectives [59, 60]. Meanwhile, surrogate approaches also offer global or local explanations by evaluating the importance of time series features with respect to prediction. These methods include the the generalization of Shapley value to time series [61], the formulation of model behavior consistency via self-supervised learning [62], and the usage of information-theoretic guided objectives for coherent explanations [63]. In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm [34, 35, 36, 37], which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
G.3 LLMs for Time Series Analysis
@@
The rapid development of LLMs [18, 19, 20] has begun to inspire new directions in time series research [64, 65]. Many existing methods fine-tune pre-trained LLMs on time series tasks and achieving state-of-the-art results in forecasting, classification, and beyond [66, 67, 68]. The textual data (e.g., domain instructions, metadata, and dataset summaries) are often encoded as prefix embeddings to enrich time series representations [69, 70, 71, 72, 73, 74]. These techniques also contribute to the emergence of time series foundation models [68, 75, 76, 77, 78]. An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series [79] or context-laden prompts representing domain knowledge [30, 32, 33], often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines [80, 31, 81, 82]. For instance, Wang et al. [82] construct a large news database for method development and perform text-to-text prediction by fine-tuning an LLM with dynamically selected news, which uses reflective selection logic to incorporate matched social events and thus augment numerical prediction. Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems [84], including time series understanding & question answering (e.g., trend and seasonal pattern recognition, similarity comparison, causality analysis, etc.[78, 86, 87, 83, 85, 88, 89], as well as language-guided inference for standard zero-shot time series tasks [83, 90, 85]. Compared to existing LLM-based methods, our approach provides a unique synergy between an explainable model and LLMs through iterative, grounded, and reflective interaction. It explicitly provides case-based explanations from both modalities, highlighting key time series patterns and refined textual segments that contribute to the prediction.
Appendix H Limitations
Our main focus has been on classification-based time series prediction to emphasize the explanation capabilities in real-world multi-modal applications, while we also provide a detailed regression-based prediction setting in the appendix. Extending our framework to broader regression-based tasks under the multi-modal contexts, such as long-term time series forecasting, and time series forecasting with domain shifts, remains an important direction for us to explore in the future work.
Appendix I Broader Impacts
Our paper presents advancements in explainable multi-modal time series prediction by integrating time series encoders with large language model-based agents. The broader impact of this work is multifaceted. It has the potential to support high-stakes decision-making in domains such as finance and healthcare by delivering more accurate predictions accompanied by reliable case-based explanations that lead to more robust analyses. The social impact of this work stems from its potential to provide a new paradigm for analyzing real-world multi-modal time series through the integration of emerging AI tools such as language agents. We emphasize the importance of responsible and ethical deployment of LLMs in time series analysis to ensure that such advancements lead to positive societal outcomes.

### v3

@@ Appendix E Designed Prompts for Experiments
Appendix F Reflection Reports for Text Refinement
@@
Figure 20: Reflection summary for text refinement on the Weather dataset.
Figure 21: Reflection summary for text refinement on the Finance dataset.
Figure 22: Reflection summary for text refinement on the Healthcare (Test-Positive) dataset.
Figure 23: Reflection summary for text refinement on the Healthcare (Mortality) dataset.
Appendix G TimeXL for Regression-based Prediction: A Finance Demonstration
@@
We conduct the experiments on the same Finance dataset, where the target value is the raw material stock price. The performance of all baselines and TimeXL is presented in Table 8. Consistent with the classification setting, TimeXL achieves the best results, and the multi-modal variants of state-of-the-art baselines (MM-iTransformer and MM-PatchTST) benefit from incorporating additional text data. This observation is further supported by the ablation study in Table 9. We also note that the text modality offers complementary information for LLMs, enhancing numerical price predictions. Furthermore, the identified prototypes provide contextual guidance, leading to additional performance gains. Overall, TimeXL outperforms all variants, underscoring the effectiveness of mutually augmented prediction. We also provide an iteration analysis to show the effectiveness of reflective and refinement process, as shown in Table 10. The prediction performance quickly improves and stabilizes over iterations, which underscores the alternation steps between predictions and reflective refinements.
@@
Figure 26: Multi-modal case-based reasoning example on the Finance dataset for regression tasks.
Finally, we provide a case study on the test set of the Finance dataset that demonstrates the effectiveness of TimeXL’s reasoning process for regression tasks, as shown in Figure 26. We present the top three most similar prototypes allocated for both time series and text modalities, where we have two key observations. 1) Even if the trend predictions align correctly in both original and refined texts, the prototype allocation varies regarding market sentiments. In the original texts, prototypes from the neutral class are also allocated, highlighting a mixed sentiment and potential fluctuations, consequently deteriorating numerical prediction accuracy from both the encoder and the LLM. In the refined texts, prototypes instead capture the underlying bearish market sentiment for raw materials, emphasizing decreased demand and weakened market conditions, effectively aligning bearish market sentiment with numerical predictions. 2) The allocated top similar time series prototypes also exhibit highly similar trends compared to the original input series, offering robust reference points for numerical prediction. To quantitatively assess it, we aggregate the retrieved prototypes using their similarity scores and compute the correlation between the input time series and the aggregated prototype series. The observed strong correlations underscores TimeXL’s efficacy in capturing intrinsic temporal patterns and providing predictive insights into future numerical stock values based on similar historical patterns.
Appendix H More Detailed Discussions of Related Work
@@
H.1 Multi-modal Time Series Analysis.
@@
In recent years, multi-modal time series analysis has gained significant traction in diverse domains such as finance, healthcare and environmental sciences [28, 11, 13]. Multiple approaches have been proposed to model interactions across different modalities, such as multi-modal fusion and cross-modal alignment and so on. Multi-modal fusion captures complementary information to enhance time series modeling, which are commonly implemented via addition [35] and concatenation [29] at the multiple levels. For example, Liu et al. [35] fuse predictions from both time series and text modalities via learnable linear weighted addition, further enhancing the predictive performance of state-of-the-art forecasters. In parallel, cross-modal alignment aims to capture the association between time series and other modalities for downstream tasks, where common techniques include attention mechanism [12, 30, 86, 31], gating functions [32], and contrastive learning [33, 34]. For instance, Bamford et al. [14] align time series and text in a shared latent space using deep encoders, enabling retrieval of specific time series sequences based on textual queries. In addition, Zheng et al. [33] perform causal structure learning in multi-modal time series by separating modality-invariant and modality-specific components through contrastive learning, enabling root cause analysis. Besides method developments, recent research efforts have also advanced the field through comprehensive benchmarks and studies [35, 36, 37], highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore representing time series as other modalities (e.g., image, text, table, graph) [38, 39, 40, 41, 42] or generating synthetic modalities [32, 43] to facilitate downstream tasks. Nevertheless, these studies tend to focus primarily on improving numerical accuracy. The deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
H.2 Time Series Explanation
@@
Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based explanations leverage saliency methods to highlight important features at different time steps [44, 45], where some methods explicitly incorporate temporal structures into models and objectives [46, 47]. Meanwhile, surrogate approaches also offer global or local explanations by evaluating the importance of time series features with respect to prediction. These methods include the the generalization of Shapley value to time series [48], the formulation of model behavior consistency via self-supervised learning [49], and the usage of information-theoretic guided objectives for coherent explanations [50]. In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm [25, 26, 27], which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
H.3 LLMs for Time Series Analysis
@@
The rapid development of LLMs [15, 16] has begun to inspire new directions in time series research [51, 52]. Many existing methods fine-tune pre-trained LLMs on time series tasks and achieving state-of-the-art results in forecasting, classification, and beyond [53, 54, 55]. The textual data (e.g., domain instructions, metadata, and dataset summaries) are often encoded as prefix embeddings to enrich time series representations [56, 57, 58, 59]. These techniques also contribute to the emergence of time series foundation models [55, 60, 61, 62]. An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series [63] or context-laden prompts representing domain knowledge [24], often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines [64, 65, 66]. For instance, Wang et al. [66] construct a large news database for method development and perform text-to-text prediction by fine-tuning an LLM with dynamically selected news, which uses reflective selection logic to incorporate matched social events and thus augment numerical prediction. Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems [68], including time series understanding & question answering (e.g., trend and seasonal pattern recognition, similarity comparison, causality analysis, etc.[67, 69, 70, 71, 72], as well as language-guided inference for standard zero-shot time series tasks [67, 69]. Compared to existing LLM-based methods, our approach provides a unique synergy between an explainable model and LLMs through iterative, grounded, and reflective interaction. It explicitly provides case-based explanations from both modalities, highlighting key time series patterns and refined textual segments that contribute to the prediction.